OpenStudy (anonymous):

Many famous Egyptian pyramids are rectangular pyramids. The formulas for the volume of a rectangular pyramid is V = 1/3 lwh. Solve this formula for w, the width of the pyramid

OpenStudy (anonymous):

Well, all we have to do is solve for w, using standard algebra. We just need to divide and multiply until we isolate the w on one side of the equation. V = (1/3) l w h Divide both sides of the equation by the other factors on the right hand side of the equation until you get something that looks like... ... ... ... ... = w

OpenStudy (anonymous):

So, if I were to divide both sides by h. We would get V / h = (1/3) l w